# Binary Black Hole Waveforms from High-Resolution GR-Athena++ Simulations

### v1

@@ III.4 Comparison with SXS waveforms
We now compare our waveforms with equivalent simulations from the SXS catalog [98, 99, 100]. In particular, we show the comparison against simulations SXS:BBH:0180, SXS:BBH:184, SXS:BBH:1179 and SXS:BBH:0167. While we report comparisons for these specific cases, we have also compared with a broader set of SXS waveforms, whose initial configurations match our waveforms. These simulations are chosen based on their low initial eccentricity and high resolution, making them suitable for informing waveform models [101]. For each SXS simulation, we employ the second extraction level and the highest resolution available. To compare these waveforms, we first use Eq. (8) to align them, and find $\Delta\phi_{22}$ and $\Delta t$. Then, the other modes are aligned by employing the same $\Delta t$ and appropriately rescaling $\Delta\phi_{22}$:
@@
Fig. (7) shows the comparison between the $(2,2)$ mode of our waveforms and SXS for all mass ratios considered. The visual agreement between the two waveforms is excellent, with the phase differences throughout the whole inspiral up to merger being consistently below $0.1$ rad. Similarly, the amplitude of the waveforms is in good agreement, with relative differences typically below $1\%$ and small oscillations observed during the inspiral phase due to some residual eccentricity in the GR-Athena++ simulations.
@@
Fig. (6), instead, shows the comparison between the $q=2$ GR-Athena++ and SXS simulations for the $(\ell,m)=\{(2,2),(2,1),(3,3),(3,2),(4,4),(5,5)\}$ modes. Similar plots for the other mass ratios are available in App. (F). Again, we find good agreement between the two simulations, with the phase differences at merger remaining at the sub-radian level for all modes investigated. The largest discrepancies are observed for the $(3,2)$ mode, with amplitude relative differences reaching ${\sim}20\%$. While it is difficult to ascertain the exact cause of these differences, we note that–as shown in Fig. 9 of Ref. [102]–that the $(3,2)$ mode is rather sensitive to numerical errors due to the center of mass drift of the binary. This effect is not corrected in the GR-Athena++ simulations presented here. Additionally, $\psi_{32}$ is affected by large numerical noise, which may also contribute to the observed inconsistencies. Overall, we find that while the dominant quadrupole modes agree to within a few tens of a radian to the SXS waveforms, the computation of precise higher modes is more challenging and may require further investigation, higher resolutions, or improved numerical techniques.
@@
To complement the time-domain comparison, we also quantify the discrepancy between the two waveform catalogs by computing mismatches. We employ the CE and LISA noise curves, and evaluate the mismatches between the starting frequency of the GR-Athena++ waveform $f_{\ell m}^{0}=m/2\omega_{22}/(2\pi)$ (multiplied by 1.25, to avoid artifacts due to FFTs) and $1024$ or $1$ Hz, respectively. We compute them between single modes only, as well as for waveform polarizations constructed using the $(\ell,m)=\{(2,2),(2,1),(3,3),(3,2),(4,4),(5,5)\}$ modes, fixing the inclination angle to $\iota=\pi/3$. We use Eq. (12) to compute the mismatches as the output is independent of the sky position of the source. When higher modes are included, we average the obtained mismatches over four values of effective polarization angle $u$. and four values of reference phase of the target waveform.
@@
Results for single mode mismatches are shown in Fig. (8). When the CE noise curve is employed, mismatches for all modes are always below the $10^{-3}$ threshold, with the notable exception of the $(5,5)$ mode for the $q=3$ case. A clear hierarchy can be observed, with the $(2,2)$ mode showing the best agreement (¯$\bar{\mathcal{F}}<10^{-4}$) for all systems considered, followed by the $(4,4)$, $(3,3)$, $(2,1)$ and $(5,5)$ modes. When the LISA noise curve is used results are comparable to those previously discussed, although a slight worsening of the agreement is observed at higher masses. Fig. (9), instead, shows the mismatches computed with polarizations constructed using the $(\ell,m)=\{(2,2),(2,1),(3,3),(3,2),(4,4),(5,5)\}$ modes. The initial frequency chosen for the mismatch calculation is $f_{0}=1.25\times\omega_{22}5/(4\pi)$, to ensure that we are focusing on a frequency region where all modes are present. The mismatches found always lie comfortably below the $10^{-3}$ threshold for both the CE and LISA noise curves.
@@ IV Conclusion and discussion
Had we accounted only for finite-resolution uncertainties in the GR-Athena++ simulations, we would have concluded that, for mass ratios of 1, 2, and 3, the waveforms are sufficiently accurate for application in next-generation detectors and LISA, where SNR can be as high as 1000. Unfortunately, the differences between waveforms at $\mathcal{I}_{+}$ obtained by running PITTNull for different inner radii are substantial. This is an unexpected finding, which indicates that even though the intrinsic error in the GR-Athena++ simulations is very small, the overall quality of our waveform is somewhat compromised. These are nevertheless among the most accurate puncture-based waveforms for BBH to date.
We compared our results with those from the SXS catalog and confirmed that our waveforms are in agreement with those produced by the SXS collaboration, with typical phase differences below $0.1$ radians throughout the inspiral and merger phases and maximum mismatches of the order of $10^{-3}$ for both the CE and LISA noise curves. These numbers indicate that our waveforms are in good agreement with those produced by the SXS collaboration for the considered number of cycles. This validation process is essential for establishing the credibility of our results and ensuring that they align with existing data from other established sources. Our work also provides another independent confirmation of the accuracy of the SXS waveform catalog, which is widely used for GW waveform modeling and data analysis applications. That said, the differences between our waveforms and those from the SXS collaboration are significantly larger than the uncertainties we have estimated for our data. These inconsistencies could be the result of gauge or initial data differences, or due to some other, unidentified source of systematic uncertainty in our waveforms or in the SXS catalog. Additional work, carefully cross validating GR-Athena++ with SXS’s SpEC [20] and SpECTRE [20] codes, is needed to understand the origin of these discrepancies.
@@ Appendix A Finite radius extraction method
where $A=1-2M/\bar{r}$ and $\bar{r}=R(1+M/(2R))^{2}$. We then use either a time domain integration or an FFI method (see App. (B)) to obtain GW strains at null infinity.
@@ Appendix B Fixed frequency integration
Finally, it is worth mentioning that to compute the second term of Eq. (16) we use FFI. Time-domain integration introduces nonphysical modulations in strain amplitude, and post-merger waveforms do not converge to zero as seen in Fig. (10).
@@ Appendix C FRE convergence
This section quantifies the numerical errors in the GW strain obtained using the FRE method due to resolution and radius extraction at null infinity. We use the same procedure outlined in Sec. III.2, and the results of the convergence analysis for the phase is shown in Fig. (11)
We extract the waveforms at $R=100$ and employ three distinct resolutions from the table (1). Each row in Fig. (11) corresponds to a specific mass ratio, while each column corresponds to a specific mode. Focusing on the dominante $(2,2)$ and $(4,4)$ modes, we use the method described in Sec. (II.5), and we find convergence behavior in all cases.
Similarly, we choose the highest resolution and perform the analysis for different extraction radii to determine the convergence order for the phase and amplitude of the FRE strains. The outcomes of this analysis are illustrated in Fig. (12). Across all cases, except $q=2$, we observe that strains converge with respect to the extraction radius [^2] While for $q=2$ run we see no convergence in phase with respect to the extraction radii, we observe a 2nd order convergence in amplitude..
@@
Fig. (13) summarizes the error analysis results. The term $err(N^{-p})$ represents the error associated to resolution of GR-Athena++, while $err(R^{-p})$ denotes the error arising form the extraction radius. Here, $p$ indicates the order of convergence found in each convergence study. We observe that the dominant error arises from the extraction radius, and the total error increases as we approach the merger. The largest errors are observed for the $q=1$ case, where the total error reaches ${\sim}0.1$ radians, and the relative amplitude difference becomes ${\sim}10\%$. Higher mass ratio runs show smaller errors, with $q=2$ approaching $\Delta A/A\sim 5\%$ and $\Delta\phi\sim 0.01$ radians. For the $q=3$ and 4 systems, the maximum total error is ${\sim}0.05$ radians in phase, and ${\sim}5\%$ in the relative amplitude difference.
@@ Appendix D FRE vs CCE
We investigate the discrepancies between the (2,2) modes of the CCE and FRE strains for our simulations. The left column of Fig. (14) compares the phase difference $|\Delta\phi|$ between the highest available resolution of GR-Athena++ (FRE strain with radius 100), and PITTNull (CCE strain with radius 50). Similarly, in the right column, we compare the amplitude between the FRE and CCE strains using $|\Delta A/\bar{A}|:=\left|(A_{\mathrm{CCE}}-A_{\mathrm{FRE}})/\bar{A}\right|,$ where $\bar{A}=(A_{\mathrm{CCE}}+A_{\mathrm{FRE}})/2$. The difference between the two methods, i.e., $\Delta\phi$ and $|\Delta A/\bar{A}|$, is shown by a black solid line. We observe that $\Delta\phi$ increases while $|\Delta A/\bar{A}|$ decreases as we approach merger. The maximum $\Delta\phi$ is ${\sim}0.1$ radians and is found for the $q=4$ run, and occurs at the merger; however, the other mass ratio runs have slightly smaller errors compared to $q=4$. Lastly, $|\Delta A/\bar{A}|$, among all runs, barely reaches $3\%$ relative difference, cf. [107].
Using numerical error estimates for the phase, obtained in Sec. III.2 and Sec. C, we plot the total numerical error $err=|err_{\mathrm{CCE}}|+|err_{\mathrm{FRE}}|$ in Fig. (14). We see that in $q=$ 1 and 3 case, the total numerical error is bigger than or comparable with $\Delta\phi$, and hence, we can expect that the difference in phases of the FRE and CCE strains are due to the numerical errors contained within each code, i.e., GR-Athena++ and PITTNull. However, for $q=$ 2 and 4, we observe dephasing is greater than the numerical error, therefore, it suggests that we are underestimating the error for one or both waveforms.
@@ Appendix E Fourier filtering in CCE
At the beginning of the evolution, and sometimes even at later stages, $C_{klm}$ in Eq. (1) exhibits noise. This noise, often stemming from numerical errors, manifests as high-frequency signals with small amplitudes, as depicted in Fig. (15). In an effort to mitigate these errors, Ref. [16] drops the high frequency content of the signal. In particular, Ref. [16] employs a Fourier transformation of the signal ${h}(t)$ to obtain the signal in Fourier space, $\tilde{h}(k)$, where $k=\frac{\omega T}{2\pi}$, $T$ is the total time, $\omega=\frac{d\phi(t)}{dt}$, and $\phi(t)$ is $\psi_{4}^{\ell m}$ phase. Then, they truncate the high-frequency noises by multiplying $\tilde{h}(k)$ with
@@
The comparison presented in Fig. (15) illustrates the efficacy of the filtering methodology proposed by [16]. When compared with the unfiltered data, this Fourier-based filtering technique demonstrates the capacity to attenuate a portion of the high-frequency noise.
@@
To assess the impact of noise present in the coefficients, we compare the FRE and CCE $\psi_{4}^{\ell m}$. The comparison is illustrated in Fig. (16). The results indicate that the FRE method yields smoother data. Nevertheless, as shown in App. (D), the FRE strains exhibit larger uncertainties compared to the CCE strains.
@@
It is important to acknowledge that as a result of extending the data and applying filtering, the tails of the data exhibit some wave patterns. Despite the enhancements made to the Fourier filtering technique, resulting in smoother $\psi_{4}^{\ell m}$ and $N_{4}^{\ell m}$ data, residual noise still persists. The comparison presented in Fig. (17) highlights the amplitude disparities between the original filtered data, the improved filtered data, and the unfiltered dataset.
@@
Given the challenges encountered with filtering methods, including data alterations and waveform noises’ persistence, the decision is made to forego filtering altogether. Additionally, using the FFI method on $\psi_{4}^{\ell m}$ to obtain strains implies filtering large frequency noises that appear in $\psi_{4}^{\ell m}$. As we observe in Fig. (2), the resultant strains are mostly free of noises.
@@
In order to evaluate the impact of the boundary condition world tube radius on waveforms, Fig. (18) presents the amplitude of $\psi_{4}^{\ell m}$ obtained using CCE for various radii. Upon examining the waveform at its onset, it becomes apparent that smaller extraction radii yield smoother results with reduced noise levels. However, it is important to note that although the smallest radius may offer smoother results, we do not observe a clean convergence with respect to the PITTNull resolutions. Consequently, the smallest radius that demonstrates convergence with respect to the PITTNull resolutions, namely, $R=50$, see Sec. (III.2), is selected for the purposes of this study.

### v3

@@ III.4 Comparison with SXS waveforms
We now compare our waveforms with equivalent simulations from the SXS catalog [103, 104, 105]. In particular, we show the comparison against simulations SXS:BBH:1155, SXS:BBH:1167, SXS:BBH:1179 and SXS:BBH:0167. While we report comparisons for these specific cases, we have also compared with a broader set of SXS waveforms, whose initial configurations match our waveforms. These simulations are chosen based on their low initial eccentricity and high resolution, making them suitable for informing waveform models [106]. For each SXS simulation, we employ the second extraction level and the highest resolution available. To compare these waveforms, we first use Eq. (8) to align them, and find $\Delta\phi_{22}$ and $\Delta t$. Then, the other modes are aligned by employing the same $\Delta t$ and appropriately rescaling $\Delta\phi_{22}$:
@@
Fig. 7 shows the comparison between the $(2,2)$ mode of our waveforms and SXS for all mass ratios considered. The visual agreement between the two waveforms is excellent, with the phase differences throughout the whole inspiral up to merger being consistently below $0.1$ rad. Similarly, the amplitude of the waveforms is in good agreement, with relative differences typically below $1\%$ and small oscillations observed during the inspiral phase due to some residual eccentricity in the GR-Athena++ simulations.
@@
Fig. 8, instead, shows the comparison between the $q=2$ GR-Athena++ and SXS simulations for the $(\ell,m)=\{(2,2),(2,1),(3,3),(3,2),(4,4),(5,5)\}$ modes. Similar plots for the other mass ratios are available in App. F. Again, we find good agreement between the two simulations, with the phase differences at merger remaining at the sub-radian level for all modes investigated. The largest discrepancies are observed for the $(3,2)$ mode, with amplitude relative differences reaching ${\sim}20\%$. While it is difficult to ascertain the exact cause of these differences, we note that–as shown in Fig. 9 of Ref. [107]–that the $(3,2)$ mode is rather sensitive to numerical errors due to the center of mass drift of the binary. This effect is not corrected in the GR-Athena++ simulations presented here. Additionally, $\psi_{32}$ is affected by large numerical noise, which may also contribute to the observed inconsistencies. Overall, we find that while the dominant quadrupole modes agree to within a few tens of a radian to the SXS waveforms, the computation of precise higher modes is more challenging and may require further investigation, higher resolutions, or improved numerical techniques.
@@
To complement the time-domain comparison, we also quantify the discrepancy between the two waveform catalogs by computing mismatches. We compute them between single modes only, as well as for waveform polarizations constructed using the $(\ell,m)=\{(2,2),(2,1),(3,3),(3,2),(4,4),(5,5)\}$ modes, fixing the inclination angle to $\iota=\pi/3$. We use Eq. (12) to compute the mismatches as the output is independent of the sky position of the source. When higher modes are included, we average the obtained mismatches over four values of effective polarization angle $u$. and four values of reference phase of the target waveform.
@@
Results for single mode mismatches are shown in Fig. 9. When the CE noise curve is employed, mismatches for all modes are always below the $10^{-3}$ threshold, with the notable exception of the $(5,5)$ mode for the $q=3$ case. A clear hierarchy can be observed, with the $(2,2)$ mode showing the best agreement (¯$\bar{\mathcal{F}}<10^{-4}$) for all systems considered, followed by the $(4,4)$, $(3,3)$, $(2,1)$ and $(5,5)$ modes. When the LISA noise curve is used results are comparable to those previously discussed, although a slight worsening of the agreement is observed at higher masses. Fig. 10, instead, shows the mismatches computed with polarizations constructed using the $(\ell,m)=\{(2,2),(2,1),(3,3),(3,2),(4,4),(5,5)\}$ modes. The initial frequency chosen for the mismatch calculation is $f_{0}=1.25\times\omega_{22}5/(4\pi)$, to ensure that we are focusing on a frequency region where all modes are present. The mismatches found always lie comfortably below the $10^{-3}$ threshold for both the CE and LISA noise curves.
@@ IV Conclusion and discussion
Our self-mismatch study for the $(2,2)$ mode of the CCE strains for mass ratios of 1, 2, 3, and 4, using the LISA noise curve and total masses ranging from $(10^{4},10^{6})M_{\odot}$, yields a mismatch of approximately $10^{-12}$ to $10^{-11}$ for a CCE inner radius of 50. Calculating the self-mismatch with the same settings but with inner radius of 100, we obtain a mismatch of approximately $10^{-5}$ to $10^{-4}$. Unfortunately, the differences between waveforms at $\mathcal{I}_{+}$ obtained by running PITTNull for different inner radii are substantial. This is an unexpected finding, which indicates that even though the intrinsic error in the GR-Athena++ simulations is very small, the overall quality of our waveform is somewhat compromised. These are nevertheless among the most accurate puncture-based waveforms for BBH to date. Additionally, our preliminary results using SpECTRE CCE [108] indicate that the CCE strains obtained using different world tube radii, namely, 50 and 100, show greater consistency in terms of error and mismatch. Consequently, we intend to conduct further experiments with SpECTRE CCE in future work.
We compared our results with those from the SXS catalog and confirmed that our waveforms are in agreement with those produced by the SXS collaboration, with typical phase differences below $0.1$ radians throughout the inspiral and merger phases and maximum mismatches of the order of $10^{-3}$ for both the CE and LISA noise curves. These numbers indicate that our waveforms are in good agreement with those produced by the SXS collaboration for the considered number of cycles. This validation process is essential for establishing the credibility of our results and ensuring that they align with existing data from other established sources. Our work also provides another independent confirmation of the accuracy of the SXS waveform catalog, which is widely used for GW waveform modeling and data analysis applications. That said, the differences between our waveforms and those from the SXS collaboration are significantly larger than the uncertainties we have estimated for our data. These inconsistencies could be the result of gauge or initial data differences, or due to some other, unidentified source of systematic uncertainty in our waveforms or in the SXS catalog. Additional work, carefully cross validating GR-Athena++ with SXS’s SpEC [105] and SpECTRE [20, 109] codes, is needed to understand the origin of these discrepancies.
@@ Appendix A Finite radius extraction method
where $A=1-2M/\bar{r}$ and $\bar{r}=R(1+M/(2R))^{2}$. We then use either a time domain integration or an FFI method (see App. B) to obtain GW strains at null infinity.
@@ Appendix B Fixed frequency integration
Finally, it is worth mentioning that to compute the second term of Eq. (16) we use FFI. Time-domain integration introduces nonphysical modulations in strain amplitude, and post-merger waveforms do not converge to zero as seen in Fig. 11.
@@ Appendix C FRE convergence
This section quantifies the numerical errors in the GW strain obtained using the FRE method due to resolution and radius extraction at null infinity. We use the same procedure outlined in Sec. III.2, and the results of the convergence analysis for the phase is shown in Fig. 12
We extract the waveforms at $R=100$ and employ three distinct resolutions from the table (1). Each row in Fig. 12 corresponds to a specific mass ratio, while each column corresponds to a specific mode. Focusing on the dominante $(2,2)$ and $(4,4)$ modes, we use the method described in Sec. (II.5), and we find convergence behavior in all cases. It is important to note that we anticipate a convergence order in resolution between 4 and 6 when the simulation is within the convergence regime (see Sec. II.1). As resolution increases, results should approach the lowest convergence order of 4, corresponding to the 4th order Runge-Kutta time integrator. Until this convergence rate is confirmed, the estimated order can vary between 4th and 6th order, as observed convergence can vary with resolution. However, identifying the same consistent convergence behavior across all simulations is complicated. A puncture solution of Einstein’s equations for BBH systems is not a regular solution and typically exhibits only $C^{2}$ continuity at the punctures [82]. Additionally, the use of different resolutions for runs with varying mass ratios further complicates this analysis.
Similarly, we choose the highest resolution and perform the analysis for different extraction radii to determine the convergence order for the phase and amplitude of the FRE strains. The outcomes of this analysis are illustrated in Fig. 13. Across all cases, except $q=2$, we observe that strains converge with respect to the extraction radius [^2] While for $q=2$ run we see no convergence in phase with respect to the extraction radii, we observe a 2nd order convergence in amplitude..
@@
Fig. 14 summarizes the error analysis results. The term $err(N^{-p})$ represents the error associated to resolution of GR-Athena++, while $err(R^{-p})$ denotes the error arising form the extraction radius. Here, $p$ indicates the order of convergence found in each convergence study. We observe that the dominant error arises from the extraction radius, and the total error increases as we approach the merger. The largest errors are observed for the $q=1$ case, where the total error reaches ${\sim}0.1$ radians, and the relative amplitude difference becomes ${\sim}10\%$. Higher mass ratio runs show smaller errors, with $q=2$ approaching $\Delta A/A\sim 5\%$ and $\Delta\phi\sim 0.01$ radians. For the $q=3$ and 4 systems, the maximum total error is ${\sim}0.05$ radians in phase, and ${\sim}5\%$ in the relative amplitude difference.
@@ Appendix D FRE vs CCE
We investigate the discrepancies between the (2,2) modes of the CCE and FRE strains for our simulations. The left column of Fig. 15 compares the phase difference $|\Delta\phi|$ between the highest available resolution of GR-Athena++ (FRE strain with radius 100), and PITTNull (CCE strain with radius 50). Similarly, in the right column, we compare the amplitude between the FRE and CCE strains using $|\Delta A/\bar{A}|:=\left|(A_{\mathrm{CCE}}-A_{\mathrm{FRE}})/\bar{A}\right|,$ where $\bar{A}=(A_{\mathrm{CCE}}+A_{\mathrm{FRE}})/2$. The difference between the two methods, i.e., $\Delta\phi$ and $|\Delta A/\bar{A}|$, is shown by a black solid line. We observe that $\Delta\phi$ increases while $|\Delta A/\bar{A}|$ decreases as we approach merger. The maximum $\Delta\phi$ is ${\sim}0.1$ radians and is found for the $q=4$ run, and occurs at the merger; however, the other mass ratio runs have slightly smaller errors compared to $q=4$. Lastly, $|\Delta A/\bar{A}|$, among all runs, barely reaches $3\%$ relative difference, cf. [113].
Using numerical error estimates for the phase, obtained in Sec. III.2 and Sec. C, we plot the total numerical error $err=|err_{\mathrm{CCE}}|+|err_{\mathrm{FRE}}|$ in Fig. 15. We see that in $q=$ 1 and 3 case, the total numerical error is bigger than or comparable with $\Delta\phi$, and hence, we can expect that the difference in phases of the FRE and CCE strains are due to the numerical errors contained within each code, i.e., GR-Athena++ and PITTNull. However, for $q=$ 2 and 4, we observe dephasing is greater than the numerical error, therefore, it suggests that we are underestimating the error for one or both waveforms.
@@ Appendix E Fourier filtering in CCE
At the beginning of the evolution, and sometimes even at later stages, $C_{klm}$ in Eq. (1) exhibits noise. This noise, often stemming from numerical errors, manifests as high-frequency signals with small amplitudes, as depicted in Fig. 16. In an effort to mitigate these errors, Ref. [16] drops the high frequency content of the signal. In particular, Ref. [16] employs a Fourier transformation of the signal ${h}(t)$ to obtain the signal in Fourier space, $\tilde{h}(k)$, where $k=\frac{\omega T}{2\pi}$, $T$ is the total time, $\omega=\frac{d\phi(t)}{dt}$, and $\phi(t)$ is $\psi_{4}^{\ell m}$ phase. Then, they truncate the high-frequency noises by multiplying $\tilde{h}(k)$ with
@@
The comparison presented in Fig. 16 illustrates the efficacy of the filtering methodology proposed by [16]. When compared with the unfiltered data, this Fourier-based filtering technique demonstrates the capacity to attenuate a portion of the high-frequency noise.
@@
To assess the impact of noise present in the coefficients, we compare the FRE and CCE $\psi_{4}^{\ell m}$. The comparison is illustrated in Fig. 17. The results indicate that the FRE method yields smoother data. Nevertheless, as shown in App. D, the FRE strains exhibit larger uncertainties compared to the CCE strains.
@@
It is important to acknowledge that as a result of extending the data and applying filtering, the tails of the data exhibit some wave patterns. Despite the enhancements made to the Fourier filtering technique, resulting in smoother $\psi_{4}^{\ell m}$ and $N_{4}^{\ell m}$ data, residual noise still persists. The comparison presented in Fig. 18 highlights the amplitude disparities between the original filtered data, the improved filtered data, and the unfiltered dataset.
@@
Given the challenges encountered with filtering methods, including data alterations and waveform noises’ persistence, the decision is made to forego filtering altogether. Additionally, using the FFI method on $\psi_{4}^{\ell m}$ to obtain strains implies filtering large frequency noises that appear in $\psi_{4}^{\ell m}$. As we observe in Fig. 2, the resultant strains are mostly free of noises.
@@
In order to evaluate the impact of the boundary condition world tube radius on waveforms, Fig. 19 presents the amplitude of $\psi_{4}^{\ell m}$ obtained using CCE for various radii. Upon examining the waveform at its onset, it becomes apparent that smaller extraction radii yield smoother results with reduced noise levels. However, it is important to note that although the smallest radius may offer smoother results, we do not observe a clean convergence with respect to the PITTNull resolutions. Consequently, the smallest radius that demonstrates convergence with respect to the PITTNull resolutions, namely, $R=50$, see Sec. (III.2), is selected for the purposes of this study.